# Charge radii measurements of exotic tin isotopes in the proximity of and

### v1

@@ Abstract
We report nuclear charge radii for the isotopes 104-134Sn, measured using two different collinear laser spectroscopy techniques at ISOLDE-CERN. These measurements clarify the archlike trend in charge radii along the isotopic chain and reveal an odd-even staggering that is more pronounced near the =50 and =82 shell closures. The observed local trends are well described by both nuclear density functional theory and valence space in-medium similarity renormalization group calculations. Both theories predict appreciable contributions from beyond-mean-field correlations to the charge radii of the neutron-deficient tin isotopes. The models fall short, however, of reproducing the magnitude of the known transition probabilities, highlighting the remaining challenges in achieving a unified description of both ground-state properties and collective phenomena.
@@
Introduction.– Understanding how nuclear shell structure and collective nuclear phenomena emerge from microscopic interactions between protons and neutrons is an essential step towards a unified description of the atomic nucleus. Collective nuclear properties are commonly enhanced in open-shell nuclei and significantly reduced around closed nuclear shells. With a closed proton shell at , and two predicted doubly magic isotopes, 100Sn and 132Sn, the tin isotopic chain provides a fertile testing ground for the development of theoretical ab-initio models [1, 2, 3, 4, 5, 6]. The properties of nuclei in the vicinity of these isotopes are expected to be accurately predicted by considering only a few valence nucleons. While the doubly magic nature of 132Sn has been well-established experimentally [7, 8, 9, 10, 11], significantly less data are available to understand the evolution of nuclear structure properties in the region from mid-shell down towards 100Sn. This is due to the considerable experimental challenges associated with producing and studying these neutron-deficient isotopes. Contradictory experimental evidence has been reported about the evolution of single-particle properties approaching 100Sn [12, 13, 14], and several questions remain open in connection with the robustness of the shell closures [1, 14, 15, 16, 17, 18, 19]. While the largest Gamow-Teller strength among all allowed -decays was found in 100Sn and supports a dominant single-particle behavior [1], the existence of low-lying energy levels and relatively large values, known down to 104Sn [15, 20, 21, 22], are indicative of a collective picture, in contrast with recent laser spectroscopy results on 49In [23] and 47Ag [16] isotopes.
@@
In this work, we present nuclear charge radii measurements for the tin isotopes 104-134Sn, using two collinear laser spectroscopy (CLS) techniques [24]. The neutron-deficient isotopes, down to 104Sn, were investigated with the Collinear Resonance Ionization Spectroscopy (CRIS) setup [25], while the neutron-rich isotopes, up to 134Sn, were measured using the Collinear Laser Spectroscopy (COLLAPS) experiment [9]. Our measurement campaigns thus extend the available data [26, 27, 28] by four isotopes on the neutron-deficient side 104-107Sn, while completing the earlier reported radii in [9] with the odd-even isotopes including 133Sn. (We note that Fig. 7 of Ref. [29] contains a data point for the charge radius of 133Sn. This has been a tentative estimate, not based on actual measurement, and was unfortunately not removed from the final version of the paper.) This exceptionally long dataset, spanning thirty-one isotopes, serves to reveal intricate changes in the mean-square charge radii towards the neutron shell closures, which can be explored by nuclear density functional theory (DFT) [30, 31] and ab initio valence space in-medium similarity renormalization group (VS-IMSRG) calculations [32, 33].
@@
Experiment.– Two CLS methods were employed to measure the charge radii of tin isotopes. Neutron-deficient isotopes, from 124Sn down to 104Sn, were measured at CRIS using two recently developed ionization schemes for probing the two atomic transitions (284 nm) and (281 nm) [34]. Neutron-rich isotopes, from 108Sn up to 134Sn were investigated at COLLAPS using two complementary transitions, (452.5 nm) and (286.3 nm). These transitions were chosen to maximize sensitivity to nuclear observables. Detailed information about the ion beam production and experimental technique used in the CRIS experiment, as well as the deduced isotope shifts and mean-square charge radii, can be found in the Supplemental Material. For the COLLAPS measurements, all details are described in [9, 35, 10]. The changes in the mean-square nuclear charge radii were extracted from the measured isotope shifts following the relation
@@
The atomic factors, and , were extracted from a standard King-plot analysis (as described in [9]) using muonic data as well as factors derived from elastic electron scattering [36]. Only the measured isotope shifts of the even stable isotopes were used in the King-plot analysis. Changes in charge radii extracted from the four transitions agree well with each other. The weighted averages were used to determine the nuclear charge radii of the tin isotopes based on the 124Sn reference value of fm of [36]. The results are presented in Table 1. The uncertainties in the extraction of charge radii are dominated by the atomic factor uncertainty.
@@
Theory.– The measured radii are confronted with predictions utilizing both the VS-IMSRG and DFT methods. The VS-IMSRG results employ two interactions derived from chiral effective field theory, 1.8/2.0 (EM) [37, 38] and N2LOGO(394) [39]. The 1.8/2.0 (EM) interaction was fitted to data of few-nucleon systems and globally reproduces ground-state energies up to nuclei [40]. This interaction has been shown to provide a reasonably good description of local variations of nuclear charge radii in the medium-mass region [41], while underestimating their absolute values [42]. In contrast, the N2LOGO(394) interaction includes the -isobar degree of freedom and was fitted to few-body data, selected nuclei, and properties of infinite nuclear matter. This interaction improves the simultaneous reproduction of energies and radii [39]. Further details of the VS-IMSRG calculations are provided in the Supplemental Material.
@@
The DFT-Fy(IVP) results were obtained with the HFB solver SkyAx for axially symmetric deformed configurations [43]. The computation of broken-pair states in odd- isotopes involves the breaking of time-reversal symmetry [44]. The applied Fy(IVP) functional [23] has been calibrated using the large dataset of ground-state properties of semi-magic nuclei from [45] augmented by the data on differential charge radii of the calcium [46], tin () and lead isotopic chains [23]. The calculated charge radii were obtained directly from the charge form factors containing relativistic corrections (including spin-orbit term) and contributions from nucleonic charge form factors [47]. The calibration of the functional produces statistical uncertainties from propagation of the uncertainties of the model parameters [45, 48]. Systematic uncertainties come from approximate angular momentum projection and from collective ground-state correlations stemming from low-lying quadrupole excitations [49]. The latter generally increase the radii and so introduce asymmetric uncertainties.
@@
Discussion.– In Figure 1, we present a detailed comparison between our measured charge radii and predictions using both VS-IMSRG and DFT approaches using the experimentally determined spins. Figure 1(a) reveals that both interactions used in the VS-IMSRG calculations underestimate the absolute charge radii. This underestimation is primarily attributed to an insufficient description of saturation [50]. In the description of saturation in VS-IMSRG, three-nucleon (3N) forces play an important role. These cannot be fixed by information from NN scattering but largely cancel in the differential radii that can thus serve as a crucial test for VS-IMSRG calculations. The Fy(IVP) DFT model provides an improved description of the radii trend and the absolute radii compared to its former variant Fy(,HFB) employed in [9] for even-even tin isotopes, particularly in the neutron-rich region and across the kink at the shell-closure.
@@
Finally, we study the odd-even staggering in Fig. 1(c). The addition of individual neutrons to the nucleus results in appreciable changes to its charge distribution due to pairing and deformation effects [41]. In our data, we observe an increased staggering as we approach the and shell closures. Notably, the characteristic inversion at , previously unobserved in tin, is now evident. The VS-IMSRG calculations for show improved agreement near these shell closures, although they tend to underestimate experiment in the mid-shell where a greater collectivity is expected. The Fy(IVP) calculations accurately predict the inversion across but systematically overestimate the staggering throughout the entire chain due to a slightly too large proton pairing predicted in mid-shell tin isotopes. It is to be noted that the estimated uncertainty in the radii below in Fy(IVP) is greater than the staggering magnitude; therefore, the results for these isotopes are omitted in Fig. 1(c).
@@
To gauge the impact of cross-shell excitation [51] across , we extended our 1.8/2.0(EM) VS-IMSRG calculations to a valence space based on the 80Zr core. This model space did, however, not include the orbit, which is known to be crucial for capturing the nuclear structure effects for isotopes with [52]. Therefore, calculations with a 80Zr core were performed from up to only. The two choices of core should yield identical results if operators induced during the IMSRG flow were not truncated, so the discrepancy between them gives an indication of the VS-IMSRG truncation error. As shown in Fig. 1, the charge radii of the light tin isotopes predicted in VS-IMSRG show relatively low sensitivity to the assumed core. We therefore conclude the charge radii description can be captured by using a 100Sn core, with calculations performed in the range from to . (As we discuss below, this is not the case for observables).
@@
We now focus on the even-even isotopes, which are simpler to interpret as the odd-neutron polarization is absent. The observed archlike trend in radii between shell closures is highlighted by removing the linear trend between the expected closed-shell isotopes 132Sn and 100Sn. The charge radius of 132Sn at the shell-closure is known. For 100Sn, the charge radius is estimated by fitting a parabolic function to the even-even radii down to 104Sn, following the same procedure described in Ref. [23]. In Fig. 2(a), the resulting archlike trend is shown after subtracting a bulk linear increase intersecting the charge radii at and .
@@
This archlike behavior can be compared to the known data on the reduced transition probabilities in Fig. 2(b). These values provide insight into quadrupole collectivity along the isotopic chain. While the measured radii reveal a clearly reduced trend towards , the trend for the values is less evident due to the large systematic uncertainty of the experimental values for the neutron-deficient isotopes. The values computed with Fy(IVP) were deduced from collective correlations, using the same approach as the correlation correction to the charge radii whereby the BCS approximation was used for simplicity. It is to be noted that these values represent the sum over all transitions. Since the lowest transition usually accounts for 90% of the total strength, the Fy(IVP) values in Fig. 2(b) serve as upper estimates and should be reduced by 10%. The Fy(IVP) results generally yield an archlike trend with a pronounced maximum at mid-shell and an only moderate reduction toward neutron-deficient isotopes. However, the mid-shell peak in is exaggerated in a similar fashion as the collective corrections for the radii. This is due to an unusually small neutron pairing predicted in that region and thus calls for further model developments.
@@
A similar archlike behavior of radii is obtained in the VS-IMSRG calculations, see Fig. 2(a). To explore the origin of this trend, spherical Hartree-Fock calculations were performed in the equal filling approximation [62], in which all -substates are filled equally, following the neutron orbitals in the order and . Hence, the observed trend is consistent with the charge radius following the neutron radius. The latter is mainly given by the sum of the single-particle mean-square radii weighted by their orbital occupation. The effect of configuration mixing is thus visualized by the comparison with the spherical HF and full VS-IMSRG results. The spherical HF results are not sensitive to collective effects, yet these calculations predict an overall archlike trend in radii. We conclude that – within the VS-IMSRG framework – the evolution of charge radii between the shell closures is dominated by the uniform isotropic growth of the mean field seen by the protons and is not due to many-body correlations. This result emphasizes that care should be taken when studying the correlations between quadrupole collectivity and nuclear charge radii [63, 64, 65, 66], as it is non-trivial to decouple the origin of different contributions to charge radii.
@@
Finally, we note that the VS-IMSRG approach significantly underestimates the magnitude of the values, as seen in Fig. 2(b). As shown earlier [67, 68] VS-IMSRG calculations based on spherical reference states typically underestimate absolute values due to missing particle-hole excitations which are needed to capture the full collective strength. The extended valence space using the 80Zr core shows an improvement, with a large increase for the magnitude of the predicted values with respect to the results using a 100Sn core, but still significantly underestimates the experimental values. This further highlights the importance of cross-shell excitations for describing the measured values of the neutron-deficient tin isotopes, as previously suggested by large-scale shell model calculations [3]. Evidently, the many-body dynamics driving the behavior of the charge radii in Fig. 2(a) is distinct from that driving the quadrupole collectivity in Fig. 2(b).
@@
As previously noted in the references [69, 41, 70, 71], reproducing the absolute magnitude of the charge radii and values in VS-IMSRG remains a challenge. While an accurate reproduction of absolute charge radii in ab initio calculations is linked to developing improved nuclear forces with correct saturation, the underestimation of values is related to the lack of collective effects in the spherical VS-IMSRG employed in this work.
Within the VS-IMSRG framework, we have shown that the observed archlike trend of charge radii between shell closures is dominated by an expansion of the nucleus due to a gradual filling of neutron shells. This is in contrast with the conclusions obtained for values and other observables [23], which require the inclusion of higher-order many-body correlations and/or non-spherical reference states. Anticipated charge radii and measurements reaching closer to 100Sn will be of great interest to complete our understanding of this frontier region of the nuclear chart.

### v2

@@ Abstract
We report nuclear charge radii for the isotopes 104-134Sn, measured using two different collinear laser spectroscopy techniques at ISOLDE-CERN. These measurements clarify the arch-like trend in charge radii along the isotopic chain and reveal an odd-even staggering that is more pronounced near the and shell closures. The observed local trends are well described by both nuclear density functional theory and valence space in-medium similarity renormalization group calculations. Both theories predict appreciable contributions from beyond-mean-field correlations to the charge radii of the neutron-deficient tin isotopes. The models, however, fall short of reproducing the magnitude of the known transition probabilities, highlighting the remaining challenges in achieving a unified description of both ground-state properties and collective phenomena.
@@
Introduction.– The mean-square charge radius is a fundamental property of the atomic nucleus, providing direct insight into its size and the spatial distribution of nuclear electric charge. Precise measurements of its evolution along isotopic chains offer unique insight into nuclear shell structure, collective phenomena, and the underlying forces. Understanding how single-nucleon and collective motion emerges from microscopic interactions between protons and neutrons is an essential step towards a unified description of the atomic nucleus. Collective nuclear properties are commonly enhanced in open-shell nuclei and significantly reduced around closed nuclear shells. In this context, the tin (Sn) isotopic chain, with its closed proton shell at , is of particular importance. Currently, it offers a long sequence of isotopes available for study, stretching across the major neutron shell closures at and . This unique landscape makes the tin chain a fertile testing ground for the development of theoretical ab initio models [hin12, Morris2018, togashi2018, gys19, mou21, Tichai:2023epe, Arthuis:2024mnl]. Many properties of nuclei in the vicinity of the doubly magic nuclei are expected to be accurately predicted by considering only a few valence nucleons. While the doubly magic nature of 132Sn has been well-established experimentally [jones2010magic, Rosiak2018a, Gorges2019, rodriguez2020doubly, Ver22], significantly less data are available to understand the evolution of nuclear structure properties in the region from mid-shell down towards 100Sn. This is due to the considerable experimental challenges associated with producing and studying these neutron-deficient isotopes. The study of these isotopes is crucial as contradictory experimental evidence has been reported about the evolution of single-particle properties approaching 100Sn [Se07, Dar10, Ba05], and several questions remain in connection with the robustness of the shell closures [hin12, Ba05, PhysRevLett.98.172501, Va07, reponen2021evidence, mou21, Nies2023, Ge2024]. While the largest Gamow-Teller strength among all allowed -decays was found in 100Sn and supports a dominant single-particle behavior [hin12], the existence of low-lying energy levels and relatively large values, known down to 104Sn [PhysRevLett.98.172501, Va07, back2013transition, guastalla2013, doornenbal2014], are indicative of a collective picture, in contrast with recent laser spectroscopy results on 49In [Karthein2024] and 47Ag [reponen2021evidence] isotopes.
@@
In this work, we present nuclear charge radii measurements for the tin isotopes 104-134Sn, using two collinear laser spectroscopy (CLS) techniques [Yan22]. The neutron-deficient isotopes, down to 104Sn, were investigated with the Collinear Resonance Ionization Spectroscopy (CRIS) setup [Kos20], while the neutron-rich isotopes, up to 134Sn, were measured using the Collinear Laser Spectroscopy (COLLAPS) experiment [Gorges2019]. Our measurement campaigns thus extend the available data [blanc2005, eberz1987, Anselment1986] by four isotopes on the neutron-deficient side 104-107Sn, while completing the earlier reported radii in [Gorges2019] with the odd-even isotopes including 133Sn 111We note that Fig. 7 of Ref. [Reinhard.2024] contains a data point for the charge radius of 133Sn. This has been a tentative estimate, not based on actual measurement, and was unfortunately not removed from the final version of the paper.. This exceptionally long dataset, spanning thirty-one isotopes, serves to reveal intricate changes in the mean-square charge radii towards the neutron shell closures, which can be explored by nuclear density functional theory (DFT) [Bender2003, nucleardft] and ab initio valence space in-medium similarity renormalization group (VS-IMSRG) calculations [Hergert:2015awm, Stro19ARNPS].
@@
Experiment.– Two CLS methods were employed to measure the charge radii of tin isotopes. Neutron-deficient isotopes, from 124Sn down to 104Sn, were measured at CRIS using two recently developed ionization schemes for probing the two atomic transitions (284 nm) and (281 nm) [Fredrik2020]. Neutron-rich isotopes, from 108Sn up to 134Sn were investigated at COLLAPS using two complementary transitions, (452.5 nm) and (286.3 nm). These transitions were chosen to maximize sensitivity to nuclear observables. Detailed information about the ion beam production and experimental technique used in the CRIS experiment, as well as the deduced isotope shifts and mean-square charge radii, can be found in the Supplemental Material. For the COLLAPS measurements, all details are described in [Gorges2019, Dey20, rodriguez2020doubly]. The changes in the mean-square nuclear charge radii were extracted from the measured isotope shifts following the relation
@@
The atomic factors, and , were extracted from a standard King-plot analysis (as described in Ref. [Gorges2019]) using muonic data as well as factors derived from elastic electron scattering [50-Sn]. Only the measured isotope shifts of the even stable isotopes were used in the King-plot analysis. Changes in charge radii extracted from the four transitions agree well with each other. The weighted averages were used to determine the nuclear charge radii of the tin isotopes based on the 124Sn reference value of fm of [50-Sn]. The results are presented in Table 1. The uncertainties in the extraction of charge radii are dominated by the atomic factor uncertainty.
@@
Theory.– The measured radii are confronted with predictions utilizing both the VS-IMSRG and DFT methods. The VS-IMSRG results employ two interactions derived from chiral effective field theory, 1.8/2.0 (EM) [Entem2003, Hebe11fits] and N2LOGO(394) [Jia20]. The 1.8/2.0 (EM) interaction was fitted to data of few-nucleon systems and globally reproduces ground-state energies up to nuclei [Stro21Drip]. This interaction has been shown to provide a reasonably good description of local variations of nuclear charge radii in the medium-mass region [deGroote2020], while underestimating their absolute values [Simo17SatFinNuc]. In contrast, the N2LOGO(394) interaction includes the -isobar degree of freedom and was fitted to few-body data, selected nuclei, and properties of infinite nuclear matter. This interaction improves the simultaneous reproduction of energies and radii [Jia20]. Further details of the VS-IMSRG calculations are provided in the Supplemental Material.
@@
The DFT-Fy(IVP) results were obtained with the HFB solver SkyAx for axially symmetric deformed configurations [SkyAx2021]. The computation of broken-pair states in odd- isotopes involves the breaking of time-reversal symmetry [Pototzky2010]. It has been calibrated using the large dataset of ground-state properties of semi-magic nuclei from [Klupfel2009] augmented by the data on differential charge radii of the calcium [miller2019a], tin [] and lead isotopic chains [Karthein2024]. The calculated charge radii were obtained directly from the charge form factors containing relativistic corrections (including spin-orbit term) and contributions from nucleonic charge form factors [DFTformfactors]. The calibration of the functional produces statistical uncertainties from propagation of the uncertainties of the model parameters [Klupfel2009, Dobaczewski2014]. Systematic uncertainties come from approximate angular momentum projection and from collective ground-state correlations stemming from low-lying quadrupole excitations [Kluepfel_2008]. The latter generally increase the radii and so introduce asymmetric uncertainties.
@@
Discussion.– In Figure 1, we present a detailed comparison between our measured charge radii and predictions using both VS-IMSRG and DFT approaches using the experimentally determined spins. Figure 1(a) reveals that both interactions used in the VS-IMSRG calculations underestimate the absolute charge radii. This underestimation is primarily attributed to an insufficient description of saturation [Simonis2017, Reinhard2016]. In the description of saturation in VS-IMSRG, three-nucleon (3N) forces play an important role. These cannot be fixed by information from NN scattering but largely cancel in the differential radii that can thus serve as a crucial test for VS-IMSRG calculations. Beyond , the VS-IMSRG calculations were performed with a different valence space only for the 1.8/2.0(EM) interaction. As the systematic uncertainties are different in different valence-space calculations, it is challenging to discuss the behavior at in the VS-ISMRG. Indeed, at , we observed a fm2 difference in the squared radii from the different valence-space calculations.
All modern nuclear DFT models (Skyrme, Fayans, Gogny, relativistic mean field) provide a good description of nuclear radii across the chart of isotopes from on, typically within fm. They differ in their performance concerning isotopic trends and details in a particular region. Fayans functionals are designed to be more flexible in tuning radius trends than other functionals [Reinhard2017, Gorges2019]. For other nuclear properties, their performance is parallel to other well-calibrated DFT models, see Ref. [ReinhardF2024]. The Fy(IVP) functional here provides a further improved description of the radii trend and the absolute radii compared to its former variant Fy(,HFB) employed in [Gorges2019] for even-even tin isotopes, particularly in the neutron-rich region and across the kink at the shell-closure.
@@
Finally, we study the odd-even staggering in Fig. 1(c). The addition of individual neutrons to the nucleus results in appreciable changes to its charge distribution due to pairing and deformation effects [deGroote2020]. In our data, we observe an increased staggering as we approach the and shell closures. In between these shell closures, we note that the staggering appears more irregular in comparison to that of neutron separation energies, as also seen along the copper chain [deGroote2020], which we attribute to the sensitivity of the charge radius to the occupation of different neutron orbitals with varying spatial extent and polarizability. In addition, our new measurement of the charge radius of reveals a discontinuity in the odd-even staggering at , similar to that observed at other major shell closures such as in Ca () and Pb () [Yan22]. The radius of suggests that a kink occurs at , which was previously identified for the even-even isotopes [Gorges2019].
@@
The VS-IMSRG calculations for show improved agreement near these shell closures, although they tend to underestimate the staggering in the mid-shell where a greater collectivity is expected. The Fy(IVP) calculations accurately predict the inversion across but systematically overestimate the staggering throughout the entire chain due to a slightly too large proton pairing predicted in mid-shell tin isotopes. It is to be noted that the estimated uncertainty in the radii below in Fy(IVP) is greater than the staggering magnitude; therefore, the results for these isotopes are omitted in Fig. 1(c).
@@
To gauge the impact of cross-shell excitation [Miyagi2020] across , we extended our 1.8/2.0(EM) VS-IMSRG calculations to a valence space based on the 80Zr core. This model space did, however, not include the orbit, which is known to be crucial for capturing the nuclear structure effects for isotopes with [Qi2012]. Therefore, calculations with a 80Zr core were performed from up to only. The two choices of core should yield identical results if operators induced during the IMSRG flow were not truncated, so the discrepancy between them is indicative of the VS-IMSRG truncation error. As shown in Fig. 1, the charge radii of the light tin isotopes predicted in VS-IMSRG show relatively low sensitivity to the assumed core. We therefore conclude the charge radii description can be captured by using a 100Sn core, with calculations performed in the range from to . (As we discuss below, this is not the case for observables).
@@
We now focus on the even-even isotopes, which are simpler to interpret as the odd-neutron polarization is absent. The observed archlike trend in radii between shell closures is highlighted by removing the linear trend between the expected closed-shell isotopes 132Sn and 100Sn. The charge radius of 132Sn at the shell-closure is known. For 100Sn, the charge radius is estimated by fitting a parabolic function to the even-even radii down to 104Sn, following the same procedure described in Ref. [Karthein2024]. In Fig. 2(a), the resulting archlike trend is shown after subtracting a bulk linear increase intersecting the charge radii at and .
@@
This archlike behavior can be compared to the known data on the reduced transition probabilities in Fig. 2(b). These values provide insight into quadrupole collectivity along the isotopic chain. While the measured radii reveal a clearly reduced trend towards , the trend for the values is less evident due to the large systematic uncertainty of the experimental values for the neutron-deficient isotopes. The values computed with Fy(IVP) were deduced from collective correlations, using the same approach as the correlation correction to the charge radii whereby the BCS approximation was used for simplicity [Kluepfel_2008]. It is to be noted that these values here represent the sum over all transitions. Since the lowest transition usually accounts for 90% of the total strength, the Fy(IVP) values in Fig. 2(b) serve as upper estimates and should be reduced by 10%. The Fy(IVP) results generally yield an archlike trend with a pronounced maximum at mid-shell and an only moderate reduction toward neutron-deficient isotopes. However, the mid-shell peak in is exaggerated in a similar fashion as the collective corrections for the radii. In addition, note that the DFT predictions of rates are still strongly varying among different functionals [Kluepfel_2008] and carry still rather large uncertainties, a feature which gives leeway for further improvement. A dedicated optimization to low-energy excitations is still a task for future development.
@@
A similar archlike behavior of radii is obtained in the VS-IMSRG calculations, see Fig. 2(a). To explore the origin of this trend, spherical Hartree-Fock calculations were performed in the equal filling approximation [Stro17], in which all -substates are filled equally, following the neutron orbitals in the order and . Hence, the observed trend is consistent with the charge radius following the neutron radius. The latter is mainly given by the sum of the single-particle mean-square radii weighted by their orbital occupation. The effect of configuration mixing is thus visualized by the comparison with the spherical HF and full VS-IMSRG results. The spherical HF results are not sensitive to collective effects, yet these calculations predict an overall archlike trend in radii. We conclude that – within the VS-IMSRG framework – the evolution of charge radii between the shell closures is dominated by the uniform isotropic growth of the mean field seen by the protons and is not due to many-body correlations. This result emphasizes that care should be taken when studying the correlations between quadrupole collectivity and nuclear charge radii [sun2017correlating, caprio2022robust, otsuka2022moments, bro22], as it is non-trivial to decouple the origin of different contributions to charge radii.
@@
Finally, we note that the VS-IMSRG approach significantly underestimates the magnitude of the values, as seen in Fig. 2(b). As shown earlier [Hend18E2, Stro22E2] VS-IMSRG calculations based on spherical reference states typically underestimate absolute values due to missing particle-hole excitations which are needed to capture the full collective strength. The extended valence space using the 80Zr core shows an improvement, with a large increase for the magnitude of the predicted values with respect to the results using a 100Sn core, but still significantly underestimates the experimental values. This further highlights the importance of cross-shell excitations for describing the measured values of the neutron-deficient tin isotopes, as previously suggested by large-scale shell model calculations [togashi2018]. On the other hand, the charge radii are not affected by much as shown in panel (a) of Fig. 1. Evidently, the many-body dynamics driving the behavior of the charge radii in Fig. 2(a) is distinct from that driving in Fig. 2(b). This emphasizes the bulk and collective nature of the charge radii and observables, respectively.
@@
As previously noted in Refs. [Parzuchowski2017, Hend18E2, Ver22, Stro22E2, LECHNER2023, Karthein2024], reproducing collective effects such as quadrupole moments and values in VS-IMSRG remains a challenge. While an accurate reproduction of absolute charge radii in ab initio calculations is linked to developing improved nuclear forces with correct saturation, the underestimation of values is related to the lack of collective effects in the spherical VS-IMSRG employed in this work.
Within the VS-IMSRG framework, we have shown that the observed archlike trend of charge radii between shell closures is dominated by an expansion of the nucleus due to a gradual filling of neutron shells. This is in contrast with the conclusions obtained for values and other observables [Karthein2024], which require the inclusion of higher-order many-body correlations and/or non-spherical reference states. Anticipated charge radii and measurements reaching closer to 100Sn will be of great interest to complete our understanding of this frontier region of the nuclear chart.
The data analysis used NumPy [numpy], SciPy [scipy], Pandas [pandas1, pandas2], iMinuit [iminuit1, iminuit2]., and qspec [Muller2025]. The figures were produced using Matplotlib [matplotlib].